how do i solve this math problem.
3x + 6= x+12
do i figure it out like this
3x+6=x+12
3x 3x
+6=3+12
+6=15
x=15/6
x=2.5

Answer by Alan3354(69443)

how do i solve this math problem.
3x + 6= x+12
do i figure it out like this
3x+6=x+12
3x 3x
+6=3+12 ***************
Adding 3x --> +6 = 3x + x+12
6 = 4x + 12
------------
+6=15 6 does not equal 15
-------------
x=15/6
x=2.5
---------
---------
6 = 4x + 12
Subtract 12
-6 = 4x
x = -6/4 = -3/2